The Ulmer Uranium Company is deciding whether or not it should open a strip mine whose net cost is $4.4 million. Net cash inflows are expected to be $27.7 million, all coming at the end of Year 1. The land must be returned to its natural state at a cost of $25 million, payable at the end of Year 2.

Should the project be accepted if r = 8%?
Should the project be accepted if r = 12%?

What is the project's MIRR at r = 8%?
What is the project's MIRR at r = 12%?

Calculate the two projects' NPVs.

Does the MIRR method lead to the same accept-reject decision as the NPV method?

Please show all work, formulas, and calculator inputs if used?
Answer #1

NPV @ 8% = -4.40 + 27.70/1.08 – 25/1.08^2 = -$0.19 million; reject as NPV is negative.

NPV @ 12% = $0.40 million; accept as NPV is positive.

MIRR at reinvestment and finance rate @ 8%

FV of cash inflow at (t=2) = 27.70*1.08 = 29.92

PV of cash outflow at (t=0) = 4.40 + 25/1.08^2 = 25.83

MIRR = (29.92/25.83)^(1/2) – 1 = 0.0761 = 7.61%; reject as rate of return (8%) is higher than MIRR (7.61%)

MIRR at reinvestment and finance rate @ 12% = 12.97%; accept as rate of return (12%) is lower than MIRR (12.97%)

MIRR is better measure than IRR does.

0 1 2 NPV @ 8% NPV @ 12% MIRR @ 8% MIRR @ 12%
Cash Flow -4.40 27.70 -25.00 -0.19 0.40 7.61% 12.92%